OpenStudy (anonymous):

what is 3/10 - w = 4/5 - 3/5w?

OpenStudy (rulnick):

3 - 10w = 8 - 6w -5 = 4w w=-4/5

OpenStudy (rulnick):

w = -5/4 I mean :)

OpenStudy (rulnick):

stoopid fingers, type right!